OpenStudy (anonymous):

The slope of the tangent line to a curve at any point (x, y) on the curve is x divided by y. What is the equation of the curve if (4, 1) is a point on the curve?

OpenStudy (irishboy123):

dy/dx = ?? build and solve.....

OpenStudy (anonymous):

I'm not sure how to begin

OpenStudy (irishboy123):

dy/dx = x / y separate + solve with boundary condition (4, 1) .....

OpenStudy (anonymous):

How do you do the boundary condition?

OpenStudy (irishboy123):

solve the differential equation first, and then use 4,1 to find the integration constant....

OpenStudy (anonymous):

would it be x+y=15

OpenStudy (irishboy123):

if x + y = 15, then y = -x + 15 so dy/dx = -1 it cannot be that, then. do you know how to solve the diff equation?